# Entropy-Triggered Retraining as Nonequilibrium Entropy Production in Deployed Machine Learning Systems

### v2

@@ Abstract
keywords:
Mathematical machine learning, Fokker–Planck equations, relative entropy, Lyapunov functionals, data drift, stochastic processes, nonequilibrium dynamics1 Introduction
@@
The modern theory of irreversible processes and entropy production originates in the foundational work of Onsager, which established reciprocity relations and identified entropy production as a structural feature of nonequilibrium systems [Onsager1931]. Subsequent developments formalized the role of probability currents in driving irreversibility and entropy generation in stochastic systems [Schnakenberg1976]. Within this framework, entropy production is not an incidental artifact of noise, but a consequence of persistent probability flow away from equilibrium.
@@ 2 Drift as Probability Flow
The Fokker–Planck equation provides a canonical description of probability flow and admits a rich analytical structure; see [Risken1989] for a comprehensive treatment.
@@
Nonzero probability currents correspond to nonequilibrium dynamics and provide a mechanism by which a fixed deployed model becomes misaligned with its environment. In nonequilibrium statistical physics, such currents are the fundamental drivers of irreversibility and entropy production, encoding sustained deviation from equilibrium through persistent probability flow [Schnakenberg1976]. This viewpoint motivates the use of entropy-based functionals as monitoring signals for deployment-time mismatch, even when only discrete-time empirical observations are available.
@@ 3 Mismatch Entropy and Its Dynamical Interpretation
which measures the statistical divergence between the evolving deployment-time data distribution and a fixed reference distribution associated with the deployed model (e.g., the training distribution or an updated reference after retraining). From an information-theoretic perspective, relative entropy provides a natural and asymmetric measure of distributional mismatch that is independent of labels or model outputs [CoverThomas2006].
@@
More generally, along nonequilibrium probability flows the entropy rate admits a decomposition analogous to entropy-balance relations in stochastic thermodynamics, featuring a nonnegative entropy production term driven by probability currents and a residual “housekeeping” term that captures sustained external driving [Seifert2005]. In the present work, this decomposition serves as a theoretical motivation rather than a directly estimated quantity. The full derivation is provided in Appendix A.
Remark 1 (Lyapunov and nonequilibrium interpretation).
@@ 5.1 Setup
wiki: daily pageviews for a popular article (Bitcoin), with log-transformed volume, rolling statistics, momentum, and weekly seasonality features.
@@ 6 Discussion
From a geometric viewpoint, relative entropy also admits an interpretation as a divergence on statistical manifolds, further supporting its role as a structural measure of model–data mismatch [Amari2016]. Practically, the main limitations concern estimation of mismatch entropy in high-dimensional settings and the choice of threshold. These are not unique to the proposed approach: any operational retraining policy requires selecting a tolerable performance–cost operating point. The advantage of entropy-triggered monitoring is that it connects this choice to an interpretable dynamical quantity driven by probability flow.
@@ Acknowledgements
AI Use Disclosure
The author used an artificial intelligence language model (ChatGPT, OpenAI; accessed December 29–31, 2025) as an assistive tool during manuscript preparation. The tool was used for expository editing and LaTeX formatting, limited symbolic checks of intermediate algebraic steps, and code drafting for experimental pipelines and figures. All mathematical derivations, modeling choices, experiments, and interpretations were independently derived, executed, and verified by the author.

### v3

@@ Abstract
Lennon J. Shikhman lshikhman3@gatech.edu
Georgia Institute of Technology
Florida Institute of Technology
Mathematical machine learning, Fokker–Planck equations, relative entropy, Lyapunov functionals, data drift, stochastic processes, nonequilibrium dynamics 111AI-based tools (ChatGPT, OpenAI) were used in a limited assistive role for editing, LaTeX formatting, symbolic checks, and code drafting; all scientific content and conclusions are the author’s own.
1 Introduction
@@
The modern theory of irreversible processes and entropy production originates in the foundational work of Onsager, which established reciprocity relations and identified entropy production as a structural feature of nonequilibrium systems Onsager (1931). Subsequent developments formalized the role of probability currents in driving irreversibility and entropy generation in stochastic systems Schnakenberg (1976). Within this framework, entropy production is not an incidental artifact of noise, but a consequence of persistent probability flow away from equilibrium.
@@ 2 Drift as Probability Flow
The Fokker–Planck equation provides a canonical description of probability flow and admits a rich analytical structure; see Risken (1989) for a comprehensive treatment.
@@
Nonzero probability currents correspond to nonequilibrium dynamics and provide a mechanism by which a fixed deployed model becomes misaligned with its environment. In nonequilibrium statistical physics, such currents are the fundamental drivers of irreversibility and entropy production, encoding sustained deviation from equilibrium through persistent probability flow Schnakenberg (1976). This viewpoint motivates the use of entropy-based functionals as monitoring signals for deployment-time mismatch, even when only discrete-time empirical observations are available.
@@ 3 Mismatch Entropy and Its Dynamical Interpretation
which measures the statistical divergence between the evolving deployment-time data distribution and a fixed reference distribution associated with the deployed model (e.g., the training distribution or an updated reference after retraining). From an information-theoretic perspective, relative entropy provides a natural and asymmetric measure of distributional mismatch that is independent of labels or model outputs Cover and Thomas (2006).
@@
More generally, along nonequilibrium probability flows the entropy rate admits a decomposition analogous to entropy-balance relations in stochastic thermodynamics, featuring a nonnegative entropy production term driven by probability currents and a residual “housekeeping” term that captures sustained external driving Seifert (2005). In the present work, this decomposition serves as a theoretical motivation rather than a directly estimated quantity. The full derivation is provided in Appendix A.
Remark 1 (Lyapunov and nonequilibrium interpretation)
@@ 5.1 Setup
Web traffic (Wikipedia pageviews): daily pageviews for a popular article (Bitcoin), with log-transformed volume, rolling statistics, momentum, and weekly seasonality features.
@@ 6 Discussion
From a geometric viewpoint, relative entropy also admits an interpretation as a divergence on statistical manifolds, further supporting its role as a structural measure of model–data mismatch Amari (2016). Practically, the main limitations concern estimation of mismatch entropy in high-dimensional settings and the choice of threshold. These are not unique to the proposed approach: any operational retraining policy requires selecting a tolerable performance–cost operating point. The advantage of entropy-triggered monitoring is that it connects this choice to an interpretable dynamical quantity driven by probability flow.